OpenStudy (anonymous):

2x -3 2x+1 ---- - ---- x -1 x + 1

OpenStudy (anonymous):

\[\frac{[(2x-3)*(x+1)]-[(2x+1)*(x-1)}{(x-1)*(x+1)}\]

OpenStudy (anonymous):

\[\frac{[2x^2-x-3]-[2x^2-x-1]}{x^2-1}\]\[-\frac{2}{x^2-1}\]